OpenStudy (anonymous):

Find the exact value of sin 82.5

OpenStudy (anonymous):

you can write it as cos 7.5 then use half angle formula which is 1+cos 2x=2cos^2 x , use it two times firstly x=15 deg then x=7.5 degrees

OpenStudy (anonymous):

I know what forumla to use I am just stuck on if I have to find the cos of 165

OpenStudy (anonymous):

I thought you had to muiltply it

OpenStudy (anonymous):

165?? i dont think you have to go that much far :P

OpenStudy (anonymous):

I'm so confused, oi

OpenStudy (anonymous):

still after that explanation?

OpenStudy (anonymous):

Ok... so.. you subtract 90-32.5=7.5

OpenStudy (anonymous):

82.3*

OpenStudy (anonymous):

Then you do \[\sqrt{1-\cos(2)(7.5)}\]

OpenStudy (anonymous):

all over two right?

OpenStudy (anonymous):

actually firstly you have to find 15 degrees by this formula coz we know the value of cos 2(15) then after use that formula again for 7.5

OpenStudy (anonymous):

so it will be \[\sqrt{1-\cos15}\]

OpenStudy (anonymous):

so you dint got what i said!!??

OpenStudy (anonymous):

no i don't because i don't get how you get 15 then divide again when you have to use to formula to muiltiply it.

OpenStudy (anonymous):

then divide my two

OpenStudy (anonymous):

i'll get 15 by {(1-cos 30 )^1/2}/2

OpenStudy (anonymous):

omg... where are you getting 30 now ??

OpenStudy (anonymous):

you dont the value of cos 30?

OpenStudy (anonymous):

..... I know the cos of 30 smart one I'm asking how you got it because from 90-82.5= 7.582=15

OpenStudy (anonymous):

wait for 5 mnts i'll upload a SS

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so are you confused how i wrote sin 82.5 = cos 7.5... right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Unit_circle

OpenStudy (anonymous):

tell if u get something from this

OpenStudy (anonymous):

Ok listen I don't have all day to read this. I am just donfused on how to start the problem.

OpenStudy (anonymous):

I don't want to sound feathery or anything but I need acutal help

OpenStudy (anonymous):

ok... i get it :D

OpenStudy (anonymous):

then getting to the point Transformation rules are:- sin 90-x = cos x, same as tanx and cot 90-x

OpenStudy (anonymous):

and sec and cosec also

OpenStudy (anonymous):

sin x= 1/cosec x, cos x=1/ secx, tan x=1/cotx

OpenStudy (anonymous):

well what do you say? :D

OpenStudy (anonymous):

Not helpful. I';m sorry. It doesn't help me. I need to know how to to set up the problem.